OpenStudy (thedj4jc):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? 4x - y = 13 4x + y = 13 4x + y = -13

OpenStudy (thedj4jc):

I think it's B, but when I try to do the first equation with the numbers I get 11. Please help?!?!?!

OpenStudy (anonymous):

parallel means has the same slpoe

OpenStudy (anonymous):

y2-y1=m(x2-x1)

OpenStudy (anonymous):

(x1,y1) (4,-3)

OpenStudy (anonymous):

the answer is B why are you not sure? the slope is the same you just have find the new intercept y=-4(4)+c -3=-4(4)+c c=13

OpenStudy (thedj4jc):

I wasn't sure because of the way the numbers changed, but now it makes sense. Thanks to both of ya'll!!